OpenStudy (anonymous):

ok they had given pendulum done 100 oscillation its amplitude reduced to 1/3 of initial value. what will be its amplitude if it completes 200 oscillation

OpenStudy (jamesj):

100 oscillations in what time period?

OpenStudy (anonymous):

dont solve just tell me ur vies dat how will u suppose to handle this ques plz plz but dont solve

OpenStudy (anonymous):

*view they didnt give time period so we can assume it constant??

OpenStudy (jamesj):

I still don't understand the question. Is it this: "In a given time period a pendulum completes 100 oscillations. "Then the pendulum has its amplitude decreased by 1/3 and it completes in the same time period 200 oscillations." If so, what's the question here.

OpenStudy (anonymous):

no they didnt said anything abu time period even i m asking can we assume it constant

OpenStudy (jamesj):

What is true about a pendulum is that if you increase the angle of maximum oscillation more than about 10 degrees, the period of the pendulum increases. The reason for that is in calculating the standard period of pendulum, we use a small angle approximation \[ \sin\theta \approx \theta \] which only holds for small angles \( \theta \). The wikipedia article has good animation of what happens when \( \theta \) gets large and there you can see the period increases. http://en.wikipedia.org/wiki/Pendulum#Period_of_oscillation This would explain the effect described in this set up.

OpenStudy (anonymous):

but james my ques how i m gonna solve it ? any hint or idea even i m trying hit and try but u can tell ur one too that it might like dis?

OpenStudy (anonymous):

u guys have any idea?

OpenStudy (anonymous):

it is saying that when pendulum completes its 100 oscilation its amplitude reduced to 1/3 of its inital means 2/3 is remain of apmlitude i guess and then they are asking if it completes 200 oscilation than what will be the value of amplitude

OpenStudy (anonymous):

i think amplitude decrease

OpenStudy (anonymous):

i need a value too.. :P

OpenStudy (jamesj):

Oh, I see. Precisely what is the amplitude, in angular terms. Use the first couple of terms of this expansion:

OpenStudy (anonymous):

god.. u r kidding james by showin me dat short formula rit ...

OpenStudy (anonymous):

i thinks in that i m not getting they are squaring or wat?? inside the bracket?

OpenStudy (shayaan_mustafa):

hello. i am back.

OpenStudy (anonymous):

with 100 oscillation amp 1/3 decrease so with 200 oscillation again decrease 1/3 im not sure

OpenStudy (jamesj):

You know that T decreases by a factor of a 1/2. Use that formula to find what value of theta_0 will make that happen.

OpenStudy (shayaan_mustafa):

may i help here? May i could do something?

OpenStudy (anonymous):

yea solve diss ques

OpenStudy (shayaan_mustafa):

ok wait. let me analyze.

OpenStudy (anonymous):

i wasted 15mins in this ques yar

OpenStudy (anonymous):

plz quick

OpenStudy (anonymous):

james formula is correct?

OpenStudy (anonymous):

yea but unable to aplly a i nvr heard abu dat one

OpenStudy (anonymous):

james formula abu period of oscillation,not amplitude

OpenStudy (shayaan_mustafa):

now what i understand. BUT THIS QUESTION LOOKS INCOMPLETE. look here first calculate time period using T=1/f T=1/100 T=0.01sec

OpenStudy (anonymous):

guys i need the ans 1/9 nw think ow can we get

OpenStudy (anonymous):

f=? y u put ocilation??

OpenStudy (anonymous):

idont understand u'r sentence?

OpenStudy (anonymous):

i mean the ans of this ques is 1/9 now just hit and try to get this ans

OpenStudy (anonymous):

how can you got it?

OpenStudy (anonymous):

no i dont got dat i had ans key :) and my ans is not matching dats why i ask for help..

OpenStudy (shayaan_mustafa):

now use the displacement equation, y=Asinwt where w is the frequency. y=(1/3) sin(100*0.01) y=(1/3)(0.0174) y=0.005m now use this displacement as, 0.005=Asin(200*0.01) 0.005=A(0.0348) A=0.143 :( is this wrong?

OpenStudy (shayaan_mustafa):

there is no other way to do it. so it tried my best. and result is before you. :((

OpenStudy (anonymous):

i want to cry.. :"( )............ )............................ god if this kind of ques coming on exam i m gonna dead

OpenStudy (shayaan_mustafa):

what is exact answer?

OpenStudy (shayaan_mustafa):

i am sure it is right.

OpenStudy (anonymous):

1/9 i meantion it too

OpenStudy (shayaan_mustafa):

1/9 is equals to 0.1. so it is right. may be.

OpenStudy (anonymous):

yes u r rit u got it congo.........

OpenStudy (shayaan_mustafa):

congo?

OpenStudy (anonymous):

but in ur equation asinwt here t is time not time period

OpenStudy (jamesj):

Heena. Let's try this one last time. Write out the complete question in your best English.

OpenStudy (anonymous):

lolz.. ok

OpenStudy (anonymous):

ithink shayan answer is wrong.

OpenStudy (anonymous):

when an osicllator completes 100 oscilation then its amplitude reduced to 1/3 of its initial value of amplitude.what will be its amplitude if an osicllator completes 200 oscilation

OpenStudy (jamesj):

This question is still very ambiguous. It means either * The time periods for the 100 and 200 oscillations are different and just a distraction. The amplitude is 3rd of what it is and that's the end of it. * We're talking about a pendulum and the period of oscillation changes. In which case we need to use the formula I wrote above * Something else.

OpenStudy (shayaan_mustafa):

i know. when i said it is true. it may be wrong. i just tried. although i am an electronics engineer not a theoritical physicist.

OpenStudy (anonymous):

guys had a idea leave it even i already wasted 30 mins in this ques better to leave this i ll ask my lecturer dont worry

OpenStudy (shayaan_mustafa):

yeah, good idea.

OpenStudy (anonymous):

thnQ for being here for help .. i always gave good ideas u knw gyan guru :P

OpenStudy (anonymous):

why gave me medal i can't solve your ques

OpenStudy (shayaan_mustafa):

well i am going. good luck every one. take care.

OpenStudy (anonymous):

hey medal is not for only those who solve the ques but its for those too who atleast tries to solve it ok

OpenStudy (anonymous):

and i think u deserve as u given ur precious time to had a luk a my ques and tries it to solve as a help thnk you so much

OpenStudy (anonymous):

this medal have not a value,i think however thanks for your attention sorry my english not well

OpenStudy (anonymous):

np i apreciate everyones work they did for helping others dats the main moto to give medals atleast people help others for medals.. and dnt think u dont solve dis medal is for ur work out in this ques ok..